OpenStudy (anonymous):

What is the x coordinate of the vertex? y = x2 + 4x + 4

OpenStudy (anonymous):

I haven't done this type of stuff in a long time

OpenStudy (callisto):

Completing square first y = x^2 + 4x + 4 = (x+2)^2 For the form y=(x-h)^2+k Vertex is at (h,k) Compare the general form and what you've got, what is the vertex?

OpenStudy (anonymous):

or you could just use the fact that -b/2a is the x-coordinated of the vertex. and to find the y coordinate, plug that into the equation to find y.